# Design and Analysis of Binaural Signal Matching with Arbitrary Microphone Arrays

### v2

@@ 7.5 ITD and ILD analysis
The performance measures studied to this point are based on the MSE, which may only partially represent human perception. Hence, a study with perceptually-motivated measures is presented next based on ITD and ILD. Both measures are calculated for sound fields comprised of a single plane wave with a DOA of $(\theta,\phi)$, where $\theta=90^{\circ}$ and $\phi$ is in the range $[0^{\circ},359^{\circ}]$ with $1^{\circ}$ resolution. The simulation parameters are similar to those used in previous sections, except for the filters $\mathbf{c}_{\text{MLS-BSM}}^{l,r}$ that were calculated with a cutoff frequency of 1.51.51.5\,1.5kHz in order to incorporate the phase of the corresponding binaural signals.
For ITD estimation, a cross-correlation based method, which was found to be an acceptable perceptual measure in [48], is utilized. With this method, the binaural signals were first passed thought a 1.51.51.5\,1.5kHz low-pass filter, followed by the calculation of the inter-aural cross-correlation (IACC):
@@
where $p^{l,r}(t)$ are the left and right time-domain binaural signals and $T$ is the total number of time samples. A cutoff frequency of 1.51.51.5\,1.5kHz was chosen since it corresponds to the frequency range in which phase or binaural time differences are important to spatial perception, as explained in Section 5. Following the IACC calculation, the ITD is estimated as:
@@
Figure 7 presents the estimated ITD and ITD errors as a function of azimuth, obtained with the six-microphone semi-circular array. In Fig 7(a), these measures are shown for static recording and reproduction conditions. The estimated ITDs are relatively accurate in this case, with ITD errors below the just noticeable difference (JND) threshold of $20\,\mu s$ and $100\,\mu s$ for the front and lateral directions, respectively [51, 52]. In addition, the errors are very similar between the BSM and its MagLS extension, which is expected since the ITD measures are dominated by the lower frequency range, at which both versions are designed similarly. The cases of compensating for head rotations during playback are presented for $\Delta\phi=30^{\circ}$ in Fig. 7(b) and for $\Delta\phi=60^{\circ}$ in Fig. 7(c). Generally, as the degree of the rotations increases, the ITD errors increase as well. More specifically, when $\Delta\phi=30^{\circ}$, the errors for lateral angles seem to increase for up to $200\,\mu s$, and up to $400\,\mu s$ when $\Delta\phi=60^{\circ}$. However, the increase in ITD errors is an additional indication that BSM may produce binaural signals with reduced quality in terms of source localization, when compensating for head rotations. This may be overcome by using an array with higher spatial resolution, such as a fully-circular array, and by using more microphones, but the study of such arrays is out of the scope of this paper.
@@ 8 Listening Experiment
This section presents a listening experiment that aims to subjectively analyze the quality of the BSM methods. More specifically, the original BSM and its MagLS extension will be compared when the acoustic environment is comprised of reverberant speech, using an egocentric microphone array that is mounted on a pair of glasses, and including listener head rotations during playback. For this purpose, two acoustic environments with different characteristics were simulated and are described next.
@@ 8.1 Setup
To generate the listening experiment signals, a point source was simulated inside a shoe-box room using the Multi-Channel Room Simulator (MCRoomSim) [55] in MATLAB [56]. The point source positions, room dimensions, and reverberation time for each room are described in Table 1. Note that two sets of room parameters were chosen to increase the diversity of the test conditions. For the source signals, four seconds long speech was used, taken from the TSP database [57], sampled at 48 kHz. Room #1 contains female speech and room #2 contains male speech. The directivity pattern of human voices was simulated using MCRoomSim as well, according to the gender of each speech signal. In order to create more realistic signals, a glasses-mounted microphone array is studied here, described in detail in [58], and with the array center position described in Table 1. The array, illustrated in Fig. 9 is comprised of six-microphones (four mounted on the glasses and additional two microphones located approximately at the ears positions), but only the four glasses-mounted microphones were used here. The array steering vectors were measured on a head and torso simulator in an anechoic chamber, as described in [58]. The corresponding pressure signals measured by the array microphones was calculated as described in (1) without adding noise.
@@ 8.2 Methodology
The listening experiment performed here follows the MUltiple Stimuli with Hidden Reference and Anchor (MUSHRA) protocol [59], but without a hidden anchor, since there is no standard anchor signal in this case. Similar experiments without a hidden anchor were studied in the past and were found to be useful, i.e., [60, 61]. The listening experiment was conducted in a static regime, but in order to study the quality of head-rotation compensation explicitly, the experiment includes signals that correspond to simulated head-rotation configurations at specific angles. Thus, for each room, three reproduction scenarios were generated, one for static binaural reproduction ($\Delta\phi=0^{\circ}$) and two for simulated head rotations with compensation - one for each rotation degree ($\Delta\phi=30^{\circ}$ and $60^{\circ}$). In each scenario, the reference was chosen to be an HOA signal of order $N=14$, leading to only minor order-truncation errors for speech signals [5, 35], calculated as described in (8). In addition, a first order Ambisonics (FOA) signal was calculated (also according to (8) but with $N_{p}=1$), to give more insight into the relative quality of each method. Hence, the four test signals in each scenario are the hidden reference (HOA), the BSM, its MagLS extension, and an FOA reproduction, leading to a total of 24 test signal. All test signals were peak-normalized to have the same dynamic range.
@@ 8.3 Results
It can be concluded that when head rotations are presented and should be compensated for, the BSM-MagLS method may produce binaural signals which are significantly better than those produced by the original BSM. However, as the degree of head rotation increases, and the compensated reproduction is becoming more challenging, the quality of BSM-MagLS may degrade significantly, compared to a HOA reproduction. Finally, BSM-MagLS produces binaural signals that are comparable to the quality achieved by HOA for static reproduction conditions, while the original BSM is closer to the quality of FOA reproduction.
@@
Figure 10: Estimated marginal means of the scores given by participants to each binaural reproduction method in the listening experiment, calculated over the two room IDs. Each screen is presented with a different category of head rotation $\Delta\phi$ indicated in the $x$-axis. The height of each bar indicates the estimated marginal mean, and the vertical black lines indicate the confidence interval of 95%.

### v3

@@ 7.5 ITD and ILD analysis
The performance measures studied to this point are based on the MSE, which may only partially represent human perception. Hence, a study with perceptually-motivated measures is presented next based on ITD and ILD. Both measures are calculated for sound fields comprised of a single plane wave with a DOA of $(\theta,\phi)$, where $\theta=90^{\circ}$ and $\phi$ is in the range $[0^{\circ},359^{\circ}]$ with $1^{\circ}$ resolution. The simulation parameters are similar to those used in previous sections, except for the filters $\mathbf{c}_{\text{BSM-MagLS}}^{l,r}$ that were calculated with a cutoff frequency of 1.51.51.5\,1.5kHz, without using cross-fade between BSM and BSM-MagLS around 1.5kHz, in order to incorporate the phase of the corresponding binaural signals.
For ITD estimation, a cross-correlation based method, which was found to be a valid perceptual measure in [51], is utilized. This method involves first low-pass filtering the binaural signals. A cutoff frequency of 1.5 kHz is used for the low-pass filter (LPF), instead of the 3 kHz LPF suggested in [51]. This lower cutoff frequency is appropriate because BSM-MagLS intentionally disregards phase information above 1.5 kHz. The filtering was followed by the calculation of the inter-aural cross-correlation (IACC):
@@
where $p^{l,r}(t)$ are the left and right time-domain binaural signals and $T$ is the total number of time samples. Following the IACC calculation, the ITD is estimated as:
@@
Figure 8 presents the estimated ITD (calculated using (42)) and ITD errors (calculated using (43)) as a function of azimuth, obtained with the six-microphone semi-circular array. In Fig 8(a), these measures are shown for static recording and reproduction conditions. The estimated ITDs are relatively accurate in this case, with ITD errors below the just noticeable difference (JND) threshold of $20\,\mu s$ and $100\,\mu s$ for the front and lateral directions, respectively [54, 55]. In addition, the errors are very similar between the BSM and its MagLS extension, which is expected since the ITD measures are dominated by the lower frequency range, at which both versions are designed similarly. The cases of compensating for head rotations during playback are presented for $\Delta\phi=30^{\circ}$ in Fig. 8(b) and for $\Delta\phi=60^{\circ}$ in Fig. 8(c). Generally, as the degree of the rotations increases, the ITD errors increase as well. More specifically, when $\Delta\phi=30^{\circ}$, the errors for lateral angles seem to increase for up to $200\,\mu s$, and up to $400\,\mu s$ when $\Delta\phi=60^{\circ}$. However, the increase in ITD errors is an additional indication that BSM may produce binaural signals with reduced quality in terms of source localization, when compensating for head rotations. This may be overcome by using an array with higher spatial resolution, such as a fully-circular array, and by using more microphones, but the study of such arrays is out of the scope of this paper.
@@ 8 Listening Experiment
This section presents a listening experiment that aims to subjectively analyze the quality of the BSM methods. More specifically, the original BSM and its MagLS extension will be compared when the acoustic environment is comprised of reverberant speech, using an egocentric microphone array that is mounted on a pair of glasses (different than the semi-circular array studied in Section 7, and including listener head rotations during playback. For this purpose, two acoustic environments with different characteristics were simulated and are described next.
@@ 8.1 Setup
To generate the listening experiment signals, a point source was simulated inside a shoe-box room using the Multi-Channel Room Simulator (MCRoomSim) [58] in MATLAB [59]. The point source positions, room dimensions, and reverberation time for each room are described in Table 1. Note that two sets of room parameters were chosen to increase the diversity of the test conditions. For the source signals, four seconds long speech was used, taken from the TSP database [60], sampled at 48 kHz. Room #1 contains female speech and room #2 contains male speech. The directivity pattern of human voices was simulated using MCRoomSim as well, according to the gender of each speech signal. In order to create more realistic signals, a glasses-mounted microphone array is studied here, described in detail in [61], and with the array center position described in Table 1. The array, illustrated in Fig. 10 is comprised of six-microphones (four mounted on the glasses and additional two microphones located approximately at the ears positions), but only the four glasses-mounted microphones were used here. The array steering vectors were measured on a head and torso simulator in an anechoic chamber, as described in [61]. The array signals were calculated by convolving the measured array steering vectors and the reverberant sound-field using SH representation. More specifically, the reverberant sound field at the origin was calculated by convolving the SH representation of the room impulse response with the clean speech signal. Then, this reverberant sound field density function was transformed to the space domain at the discrete directions corresponding to the measured array steering vectors. Finally, these two functions were convolved for each microphone position to generate the array recordings. Throughout this process, a maximal SH order of 14 was used, leading to only minor errors for this binaural signal [5, 38].
@@ 8.2 Methodology
The listening experiment performed here follows the MUltiple Stimuli with Hidden Reference and Anchor (MUSHRA) protocol [62]. The listening experiment was conducted in a static regime, but in order to study the quality of head-rotation compensation explicitly, the experiment includes signals that correspond to simulated head-rotation configurations at specific angles. Thus, for each room, three reproduction scenarios were generated, one for static binaural reproduction ($\Delta\phi=0^{\circ}$) and two for simulated head rotations with compensation - one for each rotation degree ($\Delta\phi=30^{\circ}$ and $60^{\circ}$). In each scenario, the reference was chosen to be an HOA signal of order $N=14$, leading to only minor order-truncation errors for speech signals [5, 38], calculated as described in (8). In addition, a first order Ambisonics (FOA) signal was calculated (also according to (8) but with $N_{p}=1$), and used as an anchor signal that represent a low-order reproduction. Hence, the four test signals in each scenario are the hidden reference (HOA), the BSM, its MagLS extension, and the anchor (FOA), leading to a total of 24 test signal. All test signals were normalized to the range of [-1.0, 1.0] prior to writing the final audio files.
@@ 8.3 Results
It can be concluded that BSM-MagLS with a cutoff frequency of 1.5 kHz may produce better binaural signals compared to the original BSM even when no head rotation compensation is required. This corresponds to the results in Fig. 5, with larger errors for BSM, although they correspond to a different array than the one studied here. In addition, when head rotations are presented and should be compensated for, the BSM-MagLS method may produce binaural signals which are significantly better than those produced by the original BSM. However, as the degree of head rotation increases, and the compensated reproduction is becoming more challenging, the quality of BSM-MagLS may degrade significantly, compared to a HOA reproduction. Finally, BSM-MagLS produces binaural signals that are comparable to the quality achieved by HOA for static reproduction conditions, while the original BSM is closer to the quality of FOA reproduction. While these results provide additional insights into the quality of BSM w.r.t to HOA and FOA reproduction, a more comprehensive experiment that includes other state-of-the-art binaural reproduction methods is beyond the scope of this paper and is proposed for future work.
@@
Figure 11: Box-plot of the score given by participants to each binaural reproduction method in the listening experiment, calculated over the two room IDs. The category of head rotation $\Delta\phi$ is indicated on the $x$-axis. The median is marked by the horizontal red line, outliers are marked by red plus signs, and the 25th and 75th percentiles are marked by the bottom and top colored edges, respectively. The minimal and maximal grades are marked by the black lines. Non-overlapping notches between two box-plots from the same category of head rotation indicate that the medians differ with 95% confidence.